When a Carnot engine is operated in reversed direc
Answer #1

Thermal efficiency = e = (Th - Tc)/Th

COP = Tc/(Th - Tc) = (1/e) - 1 = (1/0.87) - 1 = 1.15 - 1 = 0.15